Papier Company Limited a manufacturer of toilet paper and napkins has preferred stock of $1,000,000, paid-in capital of $900,000, paid-in capital in excess of par of $600,000, retained earnings of $60,000 (including the current year’s earnings), and 60,000 shares of common stock outstanding. This year, 2018, the company’s earnings available to common stockholders are $66,000.

Answer the following questions:

  1. Assuming that legal capital includes all paid-in capital calculate the maximum cash dividends that Papier can pay in cash dividends to each common stockholder? (3 marks)

  1. What effect will a cash dividend of $0.90 per share have on the firm’s balance sheet entries? (3 marks)

  1. If the firm cannot raise any new funds from external sources what will be the biggest constraint of the firm’s dividend payments? Provide reasons for your answer (3 marks)

  1. Papier Company Limited has a stock price of $60 per share and is considering a 3-for-2 stock split. Answer the following questions:
    1. What effect will the stock split have on Papier Company Limited equity accounts and per-share data? (3 marks)
    2. What change in stock price may result from the stock split? (3 marks)
    3. What is the maximum cash dividend per share that Papier can pay on common stock BEFORE and AFTER the stock split? (Assume that legal capital includes all paid-in capital) (5 marks)
    4. Explain the differences between stock split and stock dividends. (3 marks)
    5. What legal constraints might encourage the firm to choose a stock split over a stock dividend? (3 marks)
Answer #1

a). Maximum dividends which can be paid = retained earnings/number of shares = 60,000/60,000 = $1 per share

b). Total cash will decrease by $0.90*60,000 = 54,000. Retained earnings will decrease by the same amount.

c). Cash dividends can be given out of the cash which a company already has. In case, the company does not have sufficient cash, dividend payments will be affected so cash is the main constraint.

d-i). Current par value per share is 900,000/60,000 = $15 per share. After the split, it will be 15*2/3 = $10 per share.

Number of shares will increase to 60,000*3/2 = 90,000

Paid-in capital will remain at $900,000 (90,000*10)

EPS would be 66,000/90,000 = $0.73 per share

d-ii). Stock price will become 60*2/3 = $40 per share after the split.

d-iii). Maximum dividend before the split was $1 per share.

Maximum dividend after the split is 60,000/90,000 = $0.67 per share

d-iv). In both stock split and stock dividend, the number of outstanding shares increases. However, in a stock split, the par value per share is reduced whereas it remains unaffected in case of a stock dividend. There is no account change in a stock split but in a stock dividend, cash is moved from retained earnings to paid-in capital so it ensures that the company retains its cash. In case of a stock split, companies usually go for it when they want to bring down the stock price.

d-v). Stock dividends depend upon the retained earnings available with the company. There is no such requirement for stock splits so if retained earnings (or cash) is not sufficient, a company would choose to do a stock split.